# Modern Neural Networks for Small Tabular Datasets: The New Default for Field-Scale Digital Soil Mapping?

### v1

@@ 4.1 Deep learning versus classical machine learning
Our results demonstrated, that modern ANNs outperformed classical ML methods on the majority of prediction tasks in LimeSoDa. While our results align with recent tabular benchmarking studies (Ye et al., 2025b; Erickson et al., 2025), we extend their conclusions to the small-sample regime inherent to field-scale PSM, using datasets with as few as 30 samples(Table C1). These small-sample conditions were historically perceived as unfavorable for ANNs (Khaledian and Miller, 2020; McElfresh et al., 2023), making our findings particularly notable for field-scale PSM.
@@
The consistent performance of modern ANNs across both Low-Dimensional and High-Dimensional datasets directly challenges the established dominance of Random Forest and Linear Regression in pedometrics. Nonetheless, our results do not contradict previous findings that favored classical ML over ANN  (Shwartz-Ziv and Armon, 2022; Grinsztajn et al., 2022; McElfresh et al., 2023; Shmuel et al., 2024) as standard MLP still underperformed compared to classical ML algorithms in field-scale PSM. The critical distinction between our findings and earlier ANN evaluations lies in modern ANN architectures that either enhance the MLP framework through architectural innovations (TabM, RealMLP) or introduce fundamentally different mechanisms such as attention (ExcelFormer, T2G-Former), retrieval (TabR, ModernNCA), or in-context learning (TabPFN).
@@
Although ANNs outperformed classical ML on both Low-Dimensional and High-Dimensional datasets, the performance gap was notably smaller for High-Dimensional data. One key distinction in our High-Dimensional experiments was the use of PCA for dimensionality reduction, which was a straightforward solution to handle the unfavorable feature-to-sample ratios characteristic of spectroscopic data. While PCA has proven robust and superior to using full raw spectral data in LimeSoDa (Schmidinger et al., 2025), alternative dimensionality reduction strategies based on feature transformation (e.g. autoencoders (Rumelhart et al., 1986)), feature reduction (e.g., stochastic gates (Yamada et al., 2020)) or ensemble-based feature subsampling (Ye et al., 2025b) could further enhance ANN performance. However, comprehensive evaluation of such techniques for soil spectroscopy warrants dedicated investigation beyond the scope of this study. Regardless of these avenues for improvement, ANNs maintained their performance advantage and demonstrated general compatibility with dimensionality reduction techniques like PCA. This versatility across different data characteristics represents another key strength of selecting modern ANN architectures for PSM. It is a key advantage over tree-based models, which struggle with linearly transformed features such as those produced by PCA (Menze et al., 2009; Grinsztajn et al., 2022; Schmidinger et al., 2025).
@@
Our analysis extended the preliminary LimeSoDa benchmark of Schmidinger et al. (2025), which excluded ANNs due to their historically poor performance in tabular benchmarks. Their conclusion that different models excel in different scenarios remains partially valid. For example, we observed classical ML to be superior in the case of High-Dimensional datasets with extremely limited samples ($\leq$50), as shown in Figure 5. However, beyond this exception, modern ANNs demonstrated remarkable consistency across various field-scale PSM scenarios. TabPFN in particular maintained superior performance across nearly all dataset sizes and dimensionalities.
@@ 4.2 On the success of TabPFN
TabPFN, a pioneer in tabular in-context learning, consistently achieved the highest performance across our experiments, emerging as the top-ranked model for both Low-Dimensional and High-Dimensional datasets. This aligns with Hollmann et al. (2025) claims that TabPFN specifically excels on small to medium-sized datasets. Nonetheless, previous tabular benchmarks demonstrated TabPFN’s effectiveness mainly on datasets containing several hundred to thousands of samples (Table C1).
@@
The practical advantages of TabPFN extend beyond predictive performance. Unlike conventional ML approaches, TabPFN requires neither hyperparameter optimization nor dataset-specific training. These characteristics eliminate two of the most time-consuming and technically challenging aspects of ML workflows (Hollmann et al., 2025). For pedometricians, this simplicity rivals that of Random Forest, which has long been valued for its reliable default performance and ease of application. The combination of superior accuracy and operational simplicity positions TabPFN as an ideal modeling approach for field-scale PSM. We therefore recommend TabPFN as a new default modeling choice for every pedometrician.
Additionally, TabPFN’s success points toward opportunities for creating pedology-specific foundation models. Since the current model was trained on synthetic data, fine-tuning or retraining with real soil datasets could produce models that better understand the distinctive patterns in pedological data (Rubachev et al., 2025b; Thomas et al., 2024). Such soil-specific foundation models could combine in-context learning capabilities with learned representations of actual pedological relationships.
@@
While TabPFN excels in field-scale PSM contexts, we acknowledge scenarios where its performance may be limited. These include datasets with very large training sets, high-dimensional feature spaces without dimensionality reduction, and multi-class classification tasks (Hollmann et al., 2025; Rubachev et al., 2025a; Ye et al., 2025b). These constraints, however, fall outside the typical scope of field-scale PSM.
@@
Although TabPFN emerged as the best overall performer, we emphasize that the broader success of modern ANN architectures extends beyond TabPFN alone, and should not be overlooked. Retrieval-based models like TabR and ModernNCA, MLP-based TabM and RealMLP, and attention-based models such as T2G-Former and ExcelFormer all consistently outperformed both classical ML baselines and classical ANN architectures like MLP. Each architecture brings unique strengths and design principles that could prove valuable for specific PSM applications or inform the design of future soil-specific models. This diversity of successful approaches indicates that the integration of deep learning into PSM represents a broad methodological advance rather than dependence on any single model or architectural innovation.
@@ 4.3 Future considerations
Beyond predictive accuracy, uncertainty quantification remains critical for PSM (Breure et al., 2022; Barkov et al., 2024). While our benchmark focused on point predictions, models like TabPFN can naturally provide probabilistic predictions for uncertainty estimation, since they perform regression using target discretization (Hollmann et al., 2025; Barkov et al., 2024). Furthermore, other ANN-specific model-agnostic methods for uncertainty quantification can be evaluated, such as Laplace approximation (Kristiadi et al., 2021) or Monte Carlo dropout (Huang et al., 2025).
@@
These opportunities, alongside the proven performance of modern architectures, demonstrate that ANNs have matured sufficiently to not only advance field-scale PSM but also unlock the rapidly evolving deep learning methodological landscape for future advancements.
@@ 5 Conclusion
This study provides the first systematic assessment of modern ANNs for small tabular datasets from digital soil mapping. We implemented and evaluated a comprehensive range of recent state-of-the-art ANN architectures, including MLP-based models (MLP, TabM, RealMLP), attention-based transformers (AutoInt, FT-Transformer, T2G-Former, ExcelFormer, AMFormer), retrieval-based approaches (TabR, ModernNCA), and in-context learning models (TabPFN). We compared these architectures against established classical ML baselines (Random Forest, XGBoost, Linear Regression) across 31 diverse datasets from LimeSoDa. Through this multi-dataset and fully reproducible benchmark, we establish a new comprehensive benchmarking standard for PSM.
Our results demonstrate that many modern ANNs consistently outperform classical ML methods in field-scale PSM tasks, even under challenging small-sample conditions. The in-context learning model TabPFN emerged as a particularly strong method, surpassing Random Forest, Linear Regression, and PLSR, the long-standing default methods in PSM. We therefore recommend TabPFN as the new default model and baseline for field-scale PSM.
While classical ML retains advantages in specific areas, such as extremely small and high-dimensional datasets with fewer than 50 samples where PCA preprocessing is employed, its superiority diminishes rapidly as dataset size increases. Even for datasets containing only 50 to 100 samples, modern neural architectures can demonstrate clear performance advantages, challenging former views about sample size requirements for deep learning in general.
Our findings indicate that the recent wave of tabular deep learning research translates into tangible benefits for PSM. These results mark a substantial step forward for PSM, and as both deep learning and pedometrics fields continue to evolve, future advancements hold promise for further enhancing map accuracy and reliability in precision agriculture applications.
@@ Acknowledgements
Appendix A Deep ensembles ablation study
Figure A1: Impact of deep ensemble size on predictive performance, showing both absolute RMSE reduction and marginal improvements. Performance gains plateau after approximately 16 ensemble members, validating our choice of ensemble size and demonstrating that even small ensembles provide substantial benefits over single models in PSM.
To validate our choice of ensemble size, we conducted ablation studies varying the number of ensemble members from 1 to 64. Figure A1 presents these results.
@@ Appendix A Deep ensembles ablation study
Appendix B Detailed comparison of linear models
Figure B1: Average rank comparison of the four linear baselines across all 93 regression tasks. Lower rank values indicate better performance (best = 1). Panel (a) shows results for Low-Dimensional datasets, panel (b) for High-Dimensional datasets that were reduced with PCA. Error bars denote $\pm$ one standard error of the mean.
The Figure B1 summarizes how the linear baselines rank across all 93 regression tasks used in this study.
Appendix C Benchmark studies review
Table C1 summarizes the dataset properties used in prior influential tabular benchmarking studies. As shown, most benchmarks focused on datasets with significantly larger sample sizes and typically very low feature-to-sample ratios. Among these, only Ye et al. (2025b) incorporated datasets that closely resemble our High-Dimensional setting as part of a broader benchmark. We did not include the dataset properties from McElfresh et al. (2023), because we only found public available information from a subset of datasets used in that benchmark.
Table C1: Summary statistics of dataset properties from other previously mentioned benchmarks, compared to those used in our study, denoted as "ours". Most other benchmarks consist of datasets with larger sample sizes and low dimensional features.

### v2

@@ 4.1 Deep learning versus classical machine learning
Our results demonstrated, that modern ANNs outperformed classical ML methods on the majority of prediction tasks in LimeSoDa. While our results align with recent tabular benchmarking studies (Ye et al., 2025b; Erickson et al., 2025), we extend their conclusions to the small-sample regime inherent to field-scale DSM, using datasets with as few as 30 soil samples. These small-sample conditions were historically perceived as unfavorable for ANNs (Khaledian and Miller, 2020; McElfresh et al., 2023), making our findings particularly notable for field-scale DSM.
@@
The consistent performance of modern ANNs across both Low-Dimensional and High-Dimensional datasets directly challenges the established dominance of Random Forest and Linear Regression in pedometrics. Nonetheless, our results do not contradict previous findings that favored classical ML over ANN  (Shwartz-Ziv and Armon, 2022; Grinsztajn et al., 2022; McElfresh et al., 2023; Shmuel et al., 2024) as standard MLP still underperformed compared to classical ML algorithms in field-scale DSM. The critical distinction between our findings and earlier ANN evaluations lies in modern ANN architectures that either enhance the MLP framework through architectural innovations (TabM, RealMLP) or introduce fundamentally different mechanisms such as attention (ExcelFormer, T2G-Former), retrieval (TabR, ModernNCA), or in-context learning (TabPFN).
@@
Although modern ANNs outperformed classical ML on both Low-Dimensional and High-Dimensional datasets, the performance gap was notably smaller for High-Dimensional data. One key distinction in our High-Dimensional experiments was the use of PCA for dimensionality reduction, which was a straightforward solution to handle the unfavorable feature-to-sample ratios characteristic of spectroscopic data. While PCA has proven robust and superior to using full raw spectral data in LimeSoDa (Schmidinger et al., 2025), alternative dimensionality reduction strategies based on feature transformation (e.g., autoencoders (Rumelhart et al., 1986)), feature reduction (e.g., stochastic gates (Yamada et al., 2020)) or ensemble-based feature subsampling (Ye et al., 2025b) could further enhance ANN performance. However, comprehensive evaluation of such techniques for soil spectroscopy warrants dedicated investigation beyond the scope of this study. Regardless of these avenues for improvement, modern ANNs maintained their performance advantage and demonstrated general compatibility with dimensionality reduction techniques like PCA. This versatility across different data characteristics represents another key strength of selecting modern ANN architectures for pedometric modeling within DSM. It is a key advantage over tree-based models, which struggle with linearly transformed features such as those produced by PCA (Menze et al., 2009; Grinsztajn et al., 2022; Schmidinger et al., 2025).
@@
Our analysis extended the preliminary LimeSoDa benchmark of Schmidinger et al. (2025), which excluded ANNs due to their historically poor performance in tabular benchmarks. Their conclusion that different models excel in different scenarios remains partially valid. For example, we observed classical ML to be superior in the case of High-Dimensional datasets with extremely limited samples ($\leq$50 observations), as shown in Figure 5. However, beyond this exception, modern ANNs demonstrated remarkable consistency across various field-scale DSM scenarios. TabPFN in particular maintained superior performance across nearly all dataset sizes and dimensionalities.
The predictive performance improvements that modern ANNs demonstrate have immediate implications for pedometrics and DSM. DSM is fundamentally a task of model-based inference about unobserved locations, and its success depends on the ability of the pedometric model to represent the complex, often highly non-linear and cross-dependent, relationships between soil processes and the environment (McBratney et al., 2003). With increasingly diverse PSS and RS covariates, modern ANNs appear more capable of capturing these dependencies. Moreover, these performance improvements matter for pedological understanding as well. Strong pedometric models, paired with model-agnostic interpretability methods (e.g., Lundberg and Lee, 2017), have been suggested as a promising avenue for pedological knowledge discovery by generating hypotheses about soil-forming factors and mechanisms (Wadoux et al., 2020b). In this context, high predictive accuracy is a prerequisite, as it increases the likelihood that the pedometric model has captured relevant underlying structures (Breiman, 2003), making modern ANNs a necessary step towards potential new insights in soil science.
@@ 4.2 On the success of TabPFN
TabPFN, a pioneer in tabular in-context learning, consistently achieved the highest performance across our experiments, emerging as the top-ranked model for both Low-Dimensional and High-Dimensional datasets. This aligns with Hollmann et al. (2025) claims that TabPFN specifically excels on small to medium-sized datasets. Nonetheless, previous tabular benchmarks demonstrated TabPFN’s effectiveness mainly on datasets containing several hundred to thousands of observations (see Table S1 in the Supporting Information E).
@@
The practical advantages of TabPFN extend beyond predictive performance. TabPFN is highly computationally efficient, producing predictions on evaluated DSM datasets in seconds. Although its transformer-based attention mechanism incurs a computational complexity of $O(n^{2}+m^{2})$ with respect to sample size $n$ and feature count $m$, this cost is negligible given the small datasets typical of field-scale DSM. Furthermore, its memory requirements scale only linearly as $O(n\cdot m)$ (Hollmann et al., 2025). Unlike conventional ML approaches, TabPFN requires neither hyperparameter optimization nor dataset-specific training. These characteristics eliminate two of the most time-consuming and technically challenging aspects of ML workflows (Hollmann et al., 2025). This is especially significant in pedometric modeling, where the primary computational bottleneck often arises from iterative hyperparameter optimization rather than individual model training. Standard validation practices, such as nested cross-validation, further multiply the number of required training runs. TabPFN circumvents this bottleneck entirely, making it exceptionally practical for DSM applications. In fact, for pedometricians, this simplicity rivals that of Random Forest, which has long been valued for its reliable default performance and ease of application. The combination of superior accuracy and operational simplicity positions TabPFN as an ideal modeling approach for field-scale DSM. We therefore recommend TabPFN as a new default modeling choice for every pedometrician.
Additionally, TabPFN’s success points toward opportunities for creating pedology-specific foundation models. The synthetic data prior underlying TabPFN encodes general causal structures that enable broad generalization across different tabular domains. However, soils exhibit distinctive properties with unique structure and complex biogeochemical interactions that generic synthetic priors may not fully represent. Exploring soil-informed priors that embed fundamental domain knowledge and physical rules represents a promising research direction (Minasny et al., 2024) and could result in specialized tabular foundation models for pedometrics. Furthermore, fine-tuning TabPFN with real soil datasets could produce models that better understand these distinctive patterns in pedological data (Rubachev et al., 2025b; Thomas et al., 2024).
@@
Despite these strengths, TabPFN’s capabilities should not be overstated, as there are scenarios where its performance remains limited. These include datasets with very large training sets, high-dimensional feature spaces without dimensionality reduction, and multi-class classification tasks (Hollmann et al., 2025; Rubachev et al., 2025a; Ye et al., 2025b). These constraints, however, fall outside the typical scope of field-scale DSM.
@@
It is equally important to recognize that the advancements demonstrated in our study extend beyond TabPFN alone. Retrieval-based models like TabR and ModernNCA, MLP-based TabM and RealMLP, and attention-based models such as T2G-Former and ExcelFormer all consistently outperformed both classical ML baselines and classical ANN architectures like MLP. Each architecture brings unique strengths and design principles that could prove valuable for specific DSM applications or inform the design of future soil-specific models. We therefore interpret the success of TabPFN as a representative example of the broader capabilities that modern ANNs bring to DSM.
Finally, even the best-performing pedometric model remains subject to fundamental DSM constraints, and absolute predictive performance still varies substantially across the different LimeSoDa datasets (as shown in Figure S3 in the Supporting Information D). Advanced modeling techniques can reduce predictive uncertainty, but they represent only one component of the broader DSM workflow. Equally important are the representativeness of the sampling, influenced by the sampling design and sample size, and the quality of the features, affected by noise, preprocessing, and their overall relationship to the target soil property. Even the most advanced models will exhibit high uncertainty and poor extrapolation in under-sampled conditions, regardless of the choice of algorithm (Schmidinger et al., 2024; Žížala et al., 2024; Saurette et al., 2022). Likewise, the set of features must adequately represent soil characteristics to allow the model to find meaningful patterns. We therefore regard TabPFN, and modern ANNs in general, as a substantial but not standalone advancement that complements, rather than replaces, the fundamental sampling and feature engineering steps of DSM.
@@ 4.3 Future considerations
Beyond predictive accuracy, uncertainty quantification remains critical for pedometric modeling within DSM (Breure et al., 2022; Barkov et al., 2024). While our benchmark focused on point predictions, models like TabPFN can naturally provide probabilistic predictions for uncertainty estimation, since they perform regression using target discretization (Hollmann et al., 2025; Barkov et al., 2024). Furthermore, other ANN-specific model-agnostic methods for uncertainty quantification can be evaluated, such as Laplace approximation (Kristiadi et al., 2021) or Monte Carlo dropout (Huang et al., 2025).
@@
These opportunities, alongside the proven performance of modern architectures, demonstrate that ANNs have matured sufficiently to not only advance field-scale DSM but also unlock the rapidly evolving deep learning methodological landscape for future advancements.
@@ 5 Conclusion
This study provides the first systematic assessment of modern ANNs for small tabular datasets from DSM. We implemented and evaluated a comprehensive range of recent state-of-the-art ANN architectures, including MLP-based models (MLP, TabM, RealMLP), attention-based transformers (AutoInt, FT-Transformer, T2G-Former, ExcelFormer, AMFormer), retrieval-based approaches (TabR, ModernNCA), and in-context learning models (TabPFN). We compared these architectures against established classical ML baselines (Random Forest, XGBoost, Linear Regression) across 31 diverse datasets from LimeSoDa. Through this multi-dataset and fully reproducible benchmark, we establish a new comprehensive benchmarking standard for pedometrics and DSM.
Our results demonstrate that many modern ANNs consistently outperform classical ML methods in field-scale DSM tasks, even under challenging small-sample conditions. The in-context learning model TabPFN emerged as a particularly strong method, surpassing Random Forest, Linear Regression, and PLSR, the long-standing default methods in pedometric modeling within DSM. We therefore recommend TabPFN as the new default model and baseline for field-scale DSM. This recommendation is motivated not only by its superior predictive performance, but also by its operational simplicity, eliminating the need for hyperparameter optimization and dataset-specific training.
While classical ML retains advantages in specific areas, such as extremely small and high-dimensional datasets with fewer than 50 soil samples where PCA preprocessing is employed, its superiority diminishes rapidly as dataset size increases. Even for datasets containing only 50 to 100 soil samples, modern neural architectures can demonstrate clear performance advantages, challenging former views about sample size requirements for deep learning in general.
Our findings indicate that the recent wave of tabular deep learning research translates into tangible benefits for DSM. These results mark a substantial step forward for DSM, and as both deep learning and pedometrics fields continue to evolve, future advancements hold promise for further enhancing map accuracy and reliability in precision agriculture applications.
@@ Acknowledgements
Appendix A Detailed comparison of linear models
Figure S1: Average rank comparison of the four linear baselines across all 93 regression tasks. Lower rank values indicate better performance (best = 1). Panel (a) shows results for Low-Dimensional datasets, panel (b) for High-Dimensional datasets that were reduced with PCA. Error bars denote $\pm$ one standard error of the mean.
The Figure S1 summarizes how the linear baselines rank across all 93 regression tasks used in this study.
Appendix B Deep ensembles ablation study
Figure S2: Impact of deep ensemble size on predictive performance, showing both absolute RMSE reduction and marginal improvements. Performance gains plateau after approximately 16 ensemble members, validating our choice of ensemble size and demonstrating that even small ensembles provide substantial benefits over single models in DSM.
To validate our choice of ensemble size, we conducted ablation studies varying the number of ensemble members from 1 to 64. Figure S2 presents these results.
@@ Appendix B Deep ensembles ablation study
Appendix C Nested cross-validation and model optimization procedure
Algorithm S1 presents the complete nested cross-validation procedure employed in this study. The algorithm details the integration of hyperparameter optimization within inner folds, the prevention of data leakage through fold-specific scaling and dimensionality reduction, early stopping mechanisms, and the ensemble strategy for final predictions. The approach ensures that the outer test sets remain strictly independent from all model selection and training decisions.
Algorithm S1 Nested cross-validation with hyperparameter optimization and ensembling
Dataset $D$, learning algorithm AA, outer folds $K$, inner folds NN, optimization trials $T$, ensemble size $M$
for $i=1$ to $K$ do
   Split $D$ into $D_{\text{train}}^{i}$ and $D_{\text{test}}^{i}$
   for $t=1$ to $T$ do
    Sample hyperparameters $\theta_{t}$ from search space
    for $j=1$ to $N$ do
      Split $D_{\text{train}}^{i}$ into $D_{\text{train}}^{i,j}$ and $D_{\text{val}}^{i,j}$
      Calculate scaling statistics on $D_{\text{train}}^{i,j}$, scale $D_{\text{train}}^{i,j}$ and $D_{\text{val}}^{i,j}$
      if dataset is high-dimensional (spectroscopy) then
       Fit PCA on $D_{\text{train}}^{i,j}$, transform $D_{\text{train}}^{i,j}$ and $D_{\text{val}}^{i,j}$
      end if
      Train $A(\theta_{t})$ on $D_{\text{train}}^{i,j}$ with early stopping on $D_{\text{val}}^{i,j}$, set $\text{RMSE}_{t,j}$ on $D_{\text{val}}^{i,j}$
    end for
    Set $\text{RMSE}_{t}$ as average of $\text{RMSE}_{t,1},\dots,\text{RMSE}_{t,N}$
   end for
   Select best hyperparameters $\theta_{i}^{\star}$ with lowest $\text{RMSE}_{t}$
   for $m=1$ to $M$ do
    Split $D_{\text{train}}^{i}$ into $D_{\text{train}}^{i,m}$ and $D_{\text{val}}^{i,m}$
    Calculate scaling statistics on $D_{\text{train}}^{i,m}$, scale $D_{\text{train}}^{i,m}$, $D_{\text{val}}^{i,m}$, and $D_{\text{test}}^{i}$
    if dataset is high-dimensional (spectroscopy) then
      Fit PCA on $D_{\text{train}}^{i,m}$, transform $D_{\text{train}}^{i,m}$, $D_{\text{val}}^{i,m}$, and $D_{\text{test}}^{i}$
    end if
    Train $A(\theta_{i}^{\star})$ on $D_{\text{train}}^{i,m}$ with early stopping on $D_{\text{val}}^{i,m}$
    Predict $D_{\text{test}}^{i}$ and set predictions $\hat{y}_{i,m}$
   end for
   Set final predictions $\hat{y}_{i}$ as average of $\hat{y}_{i,1},\dots,\hat{y}_{i,M}$
end for
return Concatenated predictions $\hat{Y}=\{\hat{y}_{1},\dots,\hat{y}_{K}\}$
Appendix D Supplementary Results
Figure S3: Distribution of absolute $R^{2}$ scores across all regression tasks, including results for (a) Low-Dimensional datasets and (b) High-Dimensional spectroscopy datasets. Models are ordered vertically by their average rank for cross comparison. Whiskers extend to $1.5\times$ the interquartile range.
While rank-based aggregation provides a robust comparison of relative model performance across different regression tasks, it might abstract away the magnitude of differences across different models. Figure S3 illustrates the distribution of absolute $R^{2}$ values for all 93 regression tasks across two dataset groups.
The distributions show that better ranking of modern ANNs is supported by concrete improvements in predictive accuracy. For Low-Dimensional datasets, TabPFN shows a notably higher median $R^{2}$ with a tighter interquartile range when compared against Random Forest. Similarly, in High-Dimensional spectroscopy tasks, TabPFN clearly outperforms the best ranking classical baseline, Ridge Regression, achieving both a higher median and higher maximum predictive performance. Retrieval-based models (TabR, ModernNCA) in High-Dimensional tasks show median performance comparable to Ridge Regression, but additionally exhibit noticeably higher upper-bound results.
Appendix E Benchmark studies review
Table S1 summarizes the dataset properties used in prior influential tabular benchmarking studies. As shown, most benchmarks focused on datasets with significantly larger sample sizes and typically very low feature-to-sample ratios. Among these, only Ye et al. (2025b) incorporated datasets that closely resemble our High-Dimensional setting as part of a broader benchmark. We did not include the dataset properties from McElfresh et al. (2023), because we only found publicly available information from a subset of datasets used in that benchmark.
Table S1: Summary statistics of dataset properties from other previously mentioned benchmarks, compared to those used in our study, denoted as "ours". Most other benchmarks consist of datasets with larger sample sizes and low dimensional features.